OpenStudy (anonymous):

Which of these points lies on the line described by the equation below? y + 8 = -4(x + 7) A. (8, 7) B. (-7, -8) C. (-8, -7) D. (7, 8)

OpenStudy (anonymous):

@imqwerty @jagr2713 @Kainui

OpenStudy (anonymous):

@Victoriasushchik hi

OpenStudy (anonymous):

):

OpenStudy (victoriasushchik):

ohh wait i know

OpenStudy (anonymous):

(:

OpenStudy (victoriasushchik):

lol

OpenStudy (anonymous):

can u explain it?

OpenStudy (susanna):

y-y=m(x-x) so, y+8=-4(x+7) is really y-(-8)=-4(x-(-7)) -7 is the x, and -8 is the y. So, the point is (-7, -8)

OpenStudy (anonymous):

ry <3

OpenStudy (anonymous):

ty